• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kaufman and Moffit

    Log entry time 02:50:34 on June14,2004

    Entry number 125654

    This entry is a followup to: 125650

    keyword=Right Septum to 278.25 A

    Changed the Set Current on the Right Septum to 279.75... this was enough to bring the magnet current back up to 278.25 A ( ~= 92.58 * 3.0050).

    First attachment is the screenshot of the new Right Septum Controls.

    Second attachment is the screenshot of the last 30 hours of the Right Septum... including the recovery from the last trip.

    FIGURE 1

    FIGURE 2